OpenStudy (anonymous):

Assuming a triangle has one right angle, if side A=8, side B=11, what is the value of of side "C"

OpenStudy (anonymous):

i think a guy called pythagoras might know the answer to this question

OpenStudy (anonymous):

C^2=A^2+B^2 Therefore, 8^2 + 11^2 =185 So C =sqrt(185) =13.6

OpenStudy (anonymous):

13.60147050873544334502993737205738815299375546031776533244133798404437373029015785219574518216188210963140377282048...

OpenStudy (anonymous):

@jamesm, you have "the highest user rating" however you were the only d-bag to post an utterly useless, condescending answer.... Thanks so much for your uselessness.... Nicely done sir

OpenStudy (anonymous):

I might have loled a little at his answer and gave him a medal myself.

OpenStudy (anonymous):

my answer was telling you to use pythagoras' theorem. the fact that you couldn't use the information i provided doesn't make it useless...it just means you lack the ability to use apply the information to the formula it wasn't intended to be condescending. it was meant to be light-hearted/amusing.